# HPC-Driven Modeling with ML-Based Surrogates for Magnon-Photon Dynamics in Hybrid Quantum Systems

### v2

@@ 3 Discussion
We demonstrated hybrid quantum simulation using a magnon–photon testbed, yet the framework is readily extensible to other hybrid magnonic, cavity electrodynamic, and phononic systems  [Chen2018, Bossini2021, Hwang2024, Aspelmeyer2014]. Ongoing work therefore targets a generalized engine for quantum communication, storage, and computation that natively supports these additional interactions. Our current ML surrogate treats each mesh point as an independent one-dimensional time series, limiting its view of the device’s collective field behavior. We therefore aim to develop spatiotemporal models that learn the coupled evolution of two- and three-dimensional fields, preserve spatial coherence, and predict the full device response with high fidelity.
@@ 4 Methods
where $\gamma=-1.759\times 10^{11}$ C/kg is the gyromagnetic ratio and $\alpha$ is the dimensionless Gilbert damping constant ($10^{-5}$ to $10^{-3}$ in typical ferromagnets). The effective field $\bf{H}_{\rm eff}$} collects all magnetic interactions. Here we retain two terms essential for magnon-photon coupling: the time-dependent EM field ${\bf{H}}_{\mathrm{EM}}$, and the static bias ${\bf{H}}_{0}$. Because the ferromagnet is driven close to saturation, the magnetization magnitude is taken constant, $|{\bf{M}}|=M_{s}$, where $M_{s}$ is the (spatially dependent) saturation magnetization. We employ an iterative scheme that uses a trapezoidal discretization of the vector form of the LLG equation (5) described in [Yao2022], to solve for ${\bf{M}}^{n+1,r}$, where $n$ denotes the time step and $r$ is the counter index for the iterative procedure:
@@
We build our solver on AMReX [AMReX_JOSS], the DOE Exascale Computing Project’s block-structured adaptive mesh refinement (AMR) library. AMReX provides structured-grid PDE infrastructure with performance portability: the domain is decomposed into non-overlapping boxes (“grids”) that are distributed across MPI ranks, and execution follows an MPI+X model: X=OpenMP on multicore CPUs; or X=CUDA on NVIDIA GPUs. A single front-end code path uses AMReX loop abstractions, which select the appropriate back end at compile time. Each rank updates only the data it owns using triply nested cell loops. In the pure-MPI case, these are standard i-j-k loops; with MPI+OpenMP the loops are further tiled and assigned to threads; with MPI+CUDA AMReX launches device kernels that map cells to CUDA threads. AMReX manages data residency so arrays remain on the GPU whenever possible, minimizing host–device traffic; inter-rank communication and host/GPU transfers are therefore limited to ghost-cell exchanges performed a small number of times per time step. The efficiency and scalability of our code using MPI+CUDA simulations on NERSC Perlmutter systems is shown in Figure 6. Data from the simulation can be efficiently written using a user-defined number of MPI ranks, to prevent overwhelming the system with simultaneous writes. Visualization was performed with the publicly-available software package, VisIt [VisIt].
@@
We use the Long Expressive Memory (LEM) model [rusch2022iclr] as a robust approach for modeling complex temporal dynamics. LEM is specifically designed to capture multi-scale and long-term behaviors in systems characterized by intricate temporal structures. Its architecture offers a mathematical foundation for handling sequences with varying temporal scales, while ensuring gradient stability and high expressivity. EM fields, which exhibit a combination of rapid oscillations (e.g., high-frequency wave propagation) and slower dynamics (e.g., energy dissipation), present a challenging time-series modeling task. The LEM framework is well-suited for our scientific task due to the capability to effectively learn and predict diverse and interdependent behaviors within EM dynamics. The LEM model is rooted in multi-scale Ordinary Differential Equations (ODEs) that can be used to capture fast and slow dynamics. Let us consider a general formulation of a two-scale ODE system [rusch2022iclr], given as
@@
Physics-informed ML has been shown to effectively preserve physical properties in various domains [karniadakis2021physics]. Physics-informed loss function is adopted into the LEM framework to further improve the model’s ability to learn the underlying physics. We adopted the small-signal approximation (the dynamic magnetization in the bias direction is negligible, i.e. $M_{y}\approx 0$) in the LLG equation when deriving the physics loss term, which is
@@ 7 Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research Program, the Quantum Systems Accelerator Research Center, and the Microelectronics Science Research Center, under contract No. DE-AC02-05-CH11231. This work was also supported in part by previous breakthroughs obtained through the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under U.S. Department of Energy Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. The work at the University of Wisconsin-Madison was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of the Computational Materials Sciences Program (Y.Z. and J.-M.H.).
@@ 8 Author contributions
Z.Y. conceived the study, supervised the research, and led the manuscript writing. Z.Y. developed, configured, and executed the numerical model, generated the datasets, and performed post-processing. B.E. designed and implemented the machine learning model. J.S. and P.R. carried out model training and inference. J.S. organized and uploaded all data and codes to Zenodo. S.S. performed S-parameter extraction and analysis. S.T. performed the Floquet nonlinear analysis. Y.Z. and J.-M.H. performed the 1D theoretical analysis and the supplementary material writing. All authors discussed the results, contributed to writing and revising the manuscript, and approved the final version of the paper.

### v3

@@ 3 Discussion
We demonstrated hybrid quantum simulation using a magnon–photon testbed, yet the framework is readily extensible to other hybrid magnonic, cavity electrodynamic, and phononic systems  [10, 7, 22, 2]. Ongoing work therefore targets a generalized engine for quantum communication, storage, and computation that natively supports these additional interactions. Our current ML surrogate treats each mesh point as an independent one-dimensional time series, limiting its view of the device’s collective field behavior. We therefore aim to develop spatiotemporal models that learn the coupled evolution of two- and three-dimensional fields, preserve spatial coherence, and predict the full device response with high fidelity.
@@ 4 Methods
where $\gamma=-1.759\times 10^{11}$ C/kg is the gyromagnetic ratio and $\alpha$ is the dimensionless Gilbert damping constant ($10^{-5}$ to $10^{-3}$ in typical ferromagnets). The effective field $\bf{H}_{\rm eff}$} collects all magnetic interactions. Here we retain two terms essential for magnon-photon coupling: the time-dependent EM field ${\bf{H}}_{\mathrm{EM}}$, and the static bias ${\bf{H}}_{0}$. Because the ferromagnet is driven close to saturation, the magnetization magnitude is taken constant, $|{\bf{M}}|=M_{s}$, where $M_{s}$ is the (spatially dependent) saturation magnetization. We employ an iterative scheme that uses a trapezoidal discretization of the vector form of the LLG equation (5) described in [56], to solve for ${\bf{M}}^{n+1,r}$, where $n$ denotes the time step and $r$ is the counter index for the iterative procedure:
@@
We build our solver on AMReX [62], the DOE Exascale Computing Project’s block-structured adaptive mesh refinement (AMR) library. AMReX provides structured-grid PDE infrastructure with performance portability: the domain is decomposed into non-overlapping boxes (“grids”) that are distributed across MPI ranks, and execution follows an MPI+X model: X=OpenMP on multicore CPUs; or X=CUDA on NVIDIA GPUs. A single front-end code path uses AMReX loop abstractions, which select the appropriate back end at compile time. Each rank updates only the data it owns using triply nested cell loops. In the pure-MPI case, these are standard i-j-k loops; with MPI+OpenMP the loops are further tiled and assigned to threads; with MPI+CUDA AMReX launches device kernels that map cells to CUDA threads. AMReX manages data residency so arrays remain on the GPU whenever possible, minimizing host–device traffic; inter-rank communication and host/GPU transfers are therefore limited to ghost-cell exchanges performed a small number of times per time step. The efficiency and scalability of our code using MPI+CUDA simulations on NERSC Perlmutter systems is shown in Figure 6. Data from the simulation can be efficiently written using a user-defined number of MPI ranks, to prevent overwhelming the system with simultaneous writes. Visualization was performed with the publicly-available software package, VisIt [11].
@@
We use the Long Expressive Memory (LEM) model [42] as a robust approach for modeling complex temporal dynamics. LEM is specifically designed to capture multi-scale and long-term behaviors in systems characterized by intricate temporal structures. Its architecture offers a mathematical foundation for handling sequences with varying temporal scales, while ensuring gradient stability and high expressivity. EM fields, which exhibit a combination of rapid oscillations (e.g., high-frequency wave propagation) and slower dynamics (e.g., energy dissipation), present a challenging time-series modeling task. The LEM framework is well-suited for our scientific task due to the capability to effectively learn and predict diverse and interdependent behaviors within EM dynamics. The LEM model is rooted in multi-scale Ordinary Differential Equations (ODEs) that can be used to capture fast and slow dynamics. Let us consider a general formulation of a two-scale ODE system [42], given as
@@
Physics-informed ML has been shown to effectively preserve physical properties in various domains [23]. Physics-informed loss function is adopted into the LEM framework to further improve the model’s ability to learn the underlying physics. We adopted the small-signal approximation (the dynamic magnetization in the bias direction is negligible, i.e. $M_{y}\approx 0$) in the LLG equation when deriving the physics loss term, which is
@@ 7 Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (Award No. DE-SCL0000121). Additional support is acknowledged from the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research Program, Microelectronics Science Research Center, under contract No. DE-AC02-05-CH11231. This work was also supported in part by previous breakthroughs obtained through the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under U.S. Department of Energy Contract No. DE-AC02-05CH11231. The work at the University of Wisconsin-Madison was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of the Computational Materials Sciences Program (Y.Z. and J.-M.H.). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231.
@@ 8 Author contributions
Z.Y. conceived the study and supervised the research. Z.Y. and Y.T. led the manuscript writing. Z.Y. and A.N. developed, configured, and executed the numerical model, generated the datasets, and performed post-processing. B.E. designed and implemented the machine learning model. J.S. and P.R. carried out model training and inference. J.S. organized and uploaded all data and codes to Zenodo. S.S. performed S-parameter extraction and analysis. S.T. performed the Floquet nonlinear analysis. Y.T. helped with the training and inference. Y.Z. and J.-M.H. performed the 1D theoretical analysis and the supplementary material writing. All authors discussed the results, contributed to writing and revising the manuscript, and approved the final version of the paper.